OpenStudy (anonymous):

Help :( Assume the y varies inversely with x and y = -3 when x = 4. Write the indirect variation equation for the relationship.

OpenStudy (johnweldon1993):

Inversely means \[y = \frac{ k }{ x }\]

OpenStudy (anonymous):

Wait is xy = -12 right??

OpenStudy (johnweldon1993):

So plug in your 'x' and 'y' and solve for 'k' \[-3 = \frac{ k }{ 4 }\] multiply both sides by 4 \[k = -12\] So now your equation would be \[y = \frac{ -12 }{ x }\] Hope that helps

OpenStudy (johnweldon1993):

And yes...what you wrote is the same as what I wrote....notice if you divide both sides of YOUR equation by x...you get what I get :)

OpenStudy (anonymous):

Cool :) Thanks :)

OpenStudy (johnweldon1993):

No problem! :)